OpenStudy (anonymous):

Will give medal for answer The graph of a line passes through the two points (-2, 1) and (2, 1). What is the equation of the line written in general form? y - 1 = 0 x - y + 1 = 0 x + y - 1= 0

OpenStudy (ivettef365):

find the slope, with formula y2-y1/x2-x1 what do you get

OpenStudy (texaschic101):

exactly...find the slope first.. slope(m) = (y2 - y1) / (x2 - x1) (-2,1) x1 = -2, y1 = 1 (2,1) x2 = 2, y2 = 1 now sub your stuff in

OpenStudy (texaschic101):

Let me start you off... m = (1 - 1) / (2 - (-2)